OpenStudy (anonymous):

SOMEONE PLEASE HELP ME!!! :( Which expression describes the rule of this sequence? 3, 6, 9, 12, 15, 18 3 + (n - 1) 3n 3 + 3 (3 x n) 3 + n

OpenStudy (amistre64):

do you recall what your 3 times tables are?

jhonyy9 (jhonyy9):

so if you see it from 3 till the 6 is differents of 3 and the same from 6 till the 9 is differents of 3 so and the same from the 9 till the 12 is differents of 3 so .... this sign that the line of n atarted with 3 increasing by 3 so than will be right 3*n because 3*1=3 3*2=6 3*3=9 3*4=12 ...